Mr. Phiri has allocated his $100 budget optimally, consuming X and Y to maximise his utility at 4 utiles. When his budget was increased from $100 to $101, his utility increased to 4.05 utiles. The price of X and Y are $10 and $20 respectively.

a) What is the marginal rate of substitution of Y for X?

b) What is the slope of the budget line when Y is on the vertical axis?

c) What is the marginal utility of X at the optimum?

d) What is the marginal utility of Y at the optimum?
Answer #1

a).Marginal rate of substitution =dx2/dx1=20/10=2 units

b).slope=x/y which is 10-0/0-20=10/0=10 units

c).Marginal utility of x at the optimum is 4.05*10 =40.5 units

d).Marginal utility of x at the optimum is 4.05*20=81 units